\begin{thebibliography}{99}


\bibitem{BCZ2017}
L. A. Bokut, Y. Chen, Z. Zhang, 
Gr\"obner--Shirshov bases method for Gelfand--Dorfman--Novikov 
algebras, 
Journal of Algebra its Applications,  16(1), 1750001, 22 pp.  (2017).

\bibitem{bremner-dotsenko}
M. R. Bremner, V. Dotsenko, \emph{Algebraic Operads An Algorithmic Companion}, 2016.


\bibitem{MetNov}
A. Dauletiyarova, K. Abdukhalikov, B. K. Sartayev, On the free metabelian Novikov and metabelian Lie-admissible algebras, Communications in Mathematics, 2025, 33(3 Special issue).

\bibitem{DauSar}
A. Dauletiyarova, B. K. Sartayev, Basis of the free noncommutative Novikov algebra, Journal of Algebra and its Applications, 2024, 2550292.

\bibitem{DotsHij}
V. Dotsenko, W. Heijltjes,
\emph{Gr\"obner bases for operads}, 
\url{http://irma.math.unistra.fr/~dotsenko/Operads.html} (2019).


\bibitem{DotIsmUmi2023}
V. Dotsenko, N. Ismailov, U. Umirbaev, Polynomial identities in Novikov algebras, Mathematische Zeitschrift, 303(3), 60, (2023).

\bibitem{DzhLofwall} 
A. S. Dzhumadil’daev, C. L\"ofwall, 
Trees, free right-symmetric algebras, free Novikov algebras and identities, 
Homology, Homotopy Appl., 4(2), 165--190 (2002).

\bibitem{GD79} 
I.~M. Gelfand, I.~Ya. Dorfman, 
Hamilton operators and associated algebraic structures, 
Functional analysis and its application, {13}, no.~4, 13--30 (1979).

\bibitem{KolMashSar}
P. Kolesnikov, F. Mashurov, B. Sartayev, On Pre-Novikov Algebras and Derived Zinbiel Variety, Symmetry, Integrability and Geometry: Methods and Applications (SIGMA), 2024, 20, 17.

\bibitem{KSO2019}
P. S. Kolesnikov, B. Sartayev, A. Orazgaliev,
Gelfand--Dorfman algebras, derived identities, 
and the Manin product of operads,
Journal of Algebra {539}, 260--284 (2019).

\bibitem{Dongprop}
P. S. Kolesnikov, B. K. Sartayev, On the Dong property for a binary quadratic operad, 2025.


\bibitem{Osborn92}
J. M. Osborn, 
Novikov algebras, Nova J. Algebra Geom.
{1}, 1--14 (1992).

\bibitem{Osborn92-1}
J. M. Osborn, 
Simple Novikov algebras with an idempotent, 
Communications in Algebra {20}(9), 2729--2753 (1992).

\bibitem{Osborn94}
J. M. Osborn, 
Infinite-dimensional Novikov algebras of characteristic~0,
Journal of Algebra {167}, 146--167 (1994).

\bibitem{Pan2022}
A.S. Panasenko, Semiprime Novikov algebras, International Journal of Algebra and Computation, 32(7), pp. 1369–1378, (2022).

\bibitem{erlagol2021}
B. Sartayev, P. Kolesnikov,
Noncommutative Novikov algebras, European Journal of Mathematics, 9(2), 35, (2023).


\bibitem{Xu1996}
X. Xu, 
On simple Novikov algebras and their irreducible modules, Journal of Algebra {185}(3), 905--934 (1996).

\bibitem{Xu2001}
X. Xu, 
Classification of simple Novikov algebras and their irreducible modules of characteristic~0,
Journal of Algebra {246}(2), 673--707 (2001).

\bibitem{Zelm}
E. I. Zel'manov,
On a class of local translation invariant Lie algebras, Sov. Math. Dokl. {35}, 216--218 (1987).


\end{thebibliography}

\documentclass[a4paper, 12pt,oneside,reqno]{amsart}
\usepackage[a4paper]{geometry}
\usepackage{amssymb,amsmath,amsthm}
\usepackage[T2A]{fontenc}
\usepackage[breaklinks=true]{hyperref}

\usepackage{xcolor}

\usepackage{bbm}
\usepackage{graphicx}

\def\Var{\mathrm{Var}}
\def\dim{\mathrm{dim}}
\def\Der{\mathrm{Der}}
\def\Nov{\mathrm{Nov}}
\def\Com{\mathrm{Com}}
\def\As{\mathrm{As}}
\def\Bi{\mathrm{Bi}}

\def\Zinb{\mathrm{Zinb}}

\def\Der {\mathop {\fam 0 Der}\nolimits}

\let\<\langle
\let\>\rangle

\newtheorem{definition}{Definition}

\newtheorem{proposition}{Proposition}
\newtheorem{theorem}{Theorem}
\newtheorem{corollary}{Corollary}

\title{Differential Novikov algebras}

\author{A. Dauletiyarova}

\address{SDU University, Kaskelen, Kazakhstan}

\email{d\_aigera95@mail.ru}

\author{B. Sartayev$^{*}$}

\address{Narxoz University, Almaty, Kazakhstan}

\email{baurjai@gmail.com}

\keywords{Novikov algebra, operad, free algebra}

\subjclass[2020]{17A30, 17A50, 16R10}

\thanks{${}^{*}$Corresponding author: Bauyrzhan Sartayev   (baurjai@gmail.com)}

\begin{document}

\maketitle

\begin{abstract}
In this paper, we consider Novikov algebra with derivation and algebra obtained from its dual operad. It turns out that the obtained dual operad has a connection with bicommutative algebras. The motivation for this work comes from the white and black Manin product of the Novikov operad with itself.
\end{abstract}

\section{Introduction}

In recent years, Novikov algebras have yielded surprising results in combinatorial algebra and combinatorics. For example, the Specht property for Novikov algebras was established in \cite{DotIsmUmi2023}. One of the main combinatorial results on the free Novikov algebras is the monomial basis of $\Nov\<X\>$, see \cite{DzhLofwall}. Recently, it was shown that every Novikov algebra can be embedded into appropriate commutative algebra with a derivation \cite{BCZ2017}. The analog of this result for noncommutative Novikov algebras is given in \cite{erlagol2021}, i.e., every noncommutative Novikov algebra can be embedded into appropriate associative algebra with derivation.

The defining identities of the variety of Novikov algebras first appeared in \cite{GD79}. This work also contains the first mention of a differential associative-commutative algebra under the operation
$$a\circ b=ad(b),$$
which satisfies the following identities:
\begin{equation}\label{right-com}
    (a\circ b)\circ c=(a\circ c)\circ b
\end{equation}
\begin{equation}\label{left-sym}
    (a\circ b)\circ c-a\circ (b\circ c)=(b\circ a)\circ c-b\circ (a\circ c).
\end{equation}

The motivation for this work comes from the white and black Manin product of Novikov operad with itself. Since $\Nov=\Nov^!$, we have
\[
\Nov\bullet\Nov=((\Nov\bullet\Nov)^!)^!=(\Nov^!\bullet^!\Nov^!)^!=(\Nov\circ\Nov)^!,
\]
where $\Nov$ is an operad derived from the variety of Novikov algebras. Since there is a one-to-one correspondence between a variety of algebras $\Var$ and the quadratic operad derived from it, we will use the same terminology for both in the future.
On one hand, the white Manin product of the Novikov operad with itself gives an operad that can be embedded into the operad $\Nov$ with a derivation. On the other hand, the black Manin product of the Novikov operad with itself defines a class of algebras with a rich algebraic structure, and both operads are dual to each other.

Let $\Der\Nov$ be the operad obtained from $\Nov\circ\Nov$, and let $\Der\Nov\<X\>$ be the free algebra of the variety $\Der\Nov$ generated by a set $X$.
One of the recent combinatorial results on $\Der\Nov\<X\>$ is the following chain of inclusions:
\begin{equation}\label{inclusion}
\Der\Nov\langle X\rangle \subset 
\Nov\langle X, \partial \rangle = \Nov \langle X^{(\omega )}\rangle 
\subset \Com\langle X^{(\omega )}, d \rangle = \Com \langle X^{(\omega , \omega )} \rangle .    
\end{equation}
Here $X^{(\omega ,\omega )} = (X^{(\omega )})^{(\omega )} = \{x^{(n,m)} \mid 
x\in X, n,m\in \mathbb Z_+ \}$, 
a variable $x^{(n,m)}$ represents $d^n\partial^m(x)$. The elements of $\Der\Nov \langle X\rangle $
are exactly those polynomials in  $\Com \langle X^{(\omega , \omega )} \rangle$
that can be presented as linear combinations of monomials 
\[
x_1^{(n_1,m_1)}\dots x_k^{(n_k,m_k)},\quad \sum_i n_i = \sum_i m_i = k+1.
\]

Let us list some results on Novikov algebras.
The basis of the free metabelian Novikov algebra was given in \cite{MetNov}. The basis of the free noncommutative Novikov algebra was constructed in \cite{DauSar}.  Results on the classification of simple Novikov algebras were presented in \cite{Xu1996,Xu2001,Zelm}. Although some structural problems of Novikov algebras were studied in works such as \cite{Osborn92,Osborn92-1,Osborn94}, further research is still being conducted on other structural aspects of these algebras. For example, in \cite{Pan2022} prime and semiprime Novikov algebras were studied.

In this paper, all algebras are defined over a field of characteristic $0$.

\section{Differential Novikov algebra}

\begin{definition}
An algebra is called $\Der\Nov$ if it satisfies the following identities:
$$(a\prec b)\prec c=(a\prec c)\prec b,$$
$$a\succ(b\succ c)+(b\succ c)\prec a=b\succ(a\succ c)+(a\succ c)\prec b,$$
$$(a\succ b)\succ c-(a\succ c)\prec b=(a\succ c)\succ b-(a\succ b)\prec c$$
and
\begin{multline*}
(a\prec b)\succ c-a\prec(c\prec b)+(a\prec c)\prec b+c\succ(a\prec b)=\\
(c\prec b)\succ a-c\prec(a\prec b)+(c\prec a)\prec b+a\succ(c\prec b).
\end{multline*}
\end{definition}
These identities appear from the white Manin product of operads $\Nov$ and $\Nov$. In \cite{KolMashSar}, it was proved that
\[
\Nov\circ\Nov=\Nov\otimes\Nov
\]
which gives
\[
\dim(\Der\Nov(n))=\binom{2n-2}{n-1}^2.
\]
In \cite{KSO2019}, it was shown that operad $\Nov\circ\Var:=\Der\Var$ can always be embedded into operad $\Var^d$ under the mapping $\tau$, where $\Var^d$ is a variety of algebras $\Var$ with a derivation $d$, $\tau(a\prec b) = a d(b)$ and $\tau(a\succ b) = d(a)b$. Indeed, if the weight criterion holds then every algebra of variety $\Der\Var$ can be embedded into an appropriate algebra of variety $\Var^d$. For example, every  noncommutative Novikov algebra $\textrm{N-}\Nov\langle X\rangle$ can be embedded into appropriate differential associative algebra under the mapping
\[
\tau:\textrm{N-}\Nov \rightarrow \As^{d}, 
\]
see \cite{erlagol2021}. Since
\[
\Nov\circ\As=\Nov\otimes\As,
\]
we obtain the dimension of the operad $\textrm{N-}\Nov$:
\[
\dim(\textrm{N-}\Nov(n))=n!\binom{2n-2}{n-1}.
\]
However, such equality on the Hadamard product does not hold for the Zinbiel operad, and an example of an algebra of $\Der\Zinb$ which cannot be embedded into any algebra of $\Zinb^d$ was given \cite{KolMashSar}.

\section{Dual operad of DerNov}

Let us calculate the dual operad of the operad $\Der\Nov$ that we denote by $\Der\Nov^!$ as given in \cite{KolMashSar}. Assume 
$e_1 = x_1\succ x_2$, $e_2 = x_2\succ x_1$,  $g_1=x_1\prec x_2$ and $g_2=x_2\prec x_1$.

Consider the dual basis 
$e_1^\vee, e_2^\vee, g_1^\vee, g_2^\vee $, and denote, 
respectively, 
$e_1^\vee = y_1\succeq y_2$, 
$g_1^\vee = y_1\preceq y_2$,
then 
$e_2^\vee = - y_2\succeq y_1$ and $g_2^\vee = - y_2\preceq y_1$.
The relations on $e_1^\vee, e_1^\vee, g_1^\vee, g_2^\vee$
are exactly those that make 
the skew-symmetric bracket 
\begin{multline*}
[y_1\otimes x_1, y_2\otimes x_2]
= e_1^\vee \otimes e_1 + e_2^\vee \otimes e_2 + g_1^\vee\otimes g_1 + g_2^\vee \otimes g_2 
=\\
(y_1\succeq y_2)\otimes (x_1\succ x_2)-(y_2\succeq y_1)\otimes (x_2\succ x_1)+
(y_1\preceq y_2)\otimes (x_1\prec x_2)-(y_2\preceq y_1)\otimes (x_2\prec x_1)
\end{multline*}
to satisfy the Jacobi identity. More explicitly, we obtain
\begin{multline*}
[[y_1\otimes x_1,y_2\otimes x_2],y_3\otimes x_3]=\\
(y_1\succeq y_2)\succeq y_3\otimes(x_1\succ x_2)\succ x_3-y_3\succeq(y_1\succeq y_2)\otimes x_3\succ (x_1\succ x_2)\\
+(y_1\succeq y_2)\preceq y_3\otimes(x_1\succ x_2)\prec x_3-y_3\preceq(y_1\succeq y_2)\otimes x_3\prec (x_1\succ x_2)\\
-(y_2\succeq y_1)\succeq y_3\otimes(x_2\succ x_1)\succ x_3+y_3\succeq(y_2\succeq y_1)\otimes x_3\succ (x_2\succ x_1)\\
-(y_2\succeq y_1)\preceq y_3\otimes(x_2\succ x_1)\prec x_3+y_3\preceq(y_2\succeq y_1)\otimes x_3\prec (x_2\succ x_1)\\
+(y_1\preceq y_2)\succeq y_3\otimes(x_1\prec x_2)\succ x_3-y_3\succeq(y_1\preceq y_2)\otimes x_3\succ (x_1\prec x_2)\\
+(y_1\preceq y_2)\preceq y_3\otimes(x_1\prec x_2)\prec x_3-y_3\preceq(y_1\preceq y_2)\otimes x_3\prec (x_1\prec x_2)\\
-(y_2\preceq y_1)\succeq y_3\otimes(x_2\prec x_1)\succ x_3+y_3\succeq(y_2\preceq y_1)\otimes x_3\succ (x_2\prec x_1)\\
-(y_2\preceq y_1)\preceq y_3\otimes(x_2\prec x_1)\prec x_3+y_3\preceq(y_2\preceq y_1)\otimes x_3\prec (x_2\prec x_1).
\end{multline*}
Rewriting the right side of the tensors to the normal form in $\Der\Nov$ and collecting these basis monomials, we obtain the defining identities of the operad $\Der\Nov^!$. Since we have the inclusions (\ref{inclusion}), we can rewrite the right side of the tensors in terms of $\Com\langle X^{(\omega,\omega)}\rangle$ as follows:
\begin{multline*}
[[y_1\otimes x_1,y_2\otimes x_2],y_3\otimes x_3]=
(y_1\succeq y_2)\succeq y_3\otimes(x_1'\circ x_2)'\circ x_3-y_3\succeq(y_1\succeq y_2)\otimes x_3'\circ (x_1'\circ x_2)\\
+(y_1\succeq y_2)\preceq y_3\otimes(x_1'\circ x_2)\circ x_3'-y_3\preceq(y_1\succeq y_2)\otimes x_3\circ (x_1'\circ x_2)'\\
-(y_2\succeq y_1)\succeq y_3\otimes(x_2'\circ x_1)'\circ x_3+y_3\succeq(y_2\succeq y_1)\otimes x_3'\circ (x_2'\circ x_1)\\
-(y_2\succeq y_1)\preceq y_3\otimes(x_2'\circ x_1)\circ x_3'+y_3\preceq(y_2\succeq y_1)\otimes x_3\circ (x_2'\circ x_1)'\\
+(y_1\preceq y_2)\succeq y_3\otimes(x_1\circ x_2')'\circ x_3-y_3\succeq(y_1\preceq y_2)\otimes x_3'\circ (x_1\circ x_2')\\
+(y_1\preceq y_2)\preceq y_3\otimes(x_1\circ x_2')\circ x_3'-y_3\preceq(y_1\preceq y_2)\otimes x_3\circ (x_1\circ x_2')'
\end{multline*}
\vspace*{-\baselineskip}
\vspace*{-\baselineskip}
\begin{multline*}
-(y_2\preceq y_1)\succeq y_3\otimes(x_2\circ x_1')'\circ x_3+y_3\succeq(y_2\preceq y_1)\otimes x_3'\circ (x_2\circ x_1')\\
-(y_2\preceq y_1)\preceq y_3\otimes(x_2\circ x_1')\circ x_3'+y_3\preceq(y_2\preceq y_1)\otimes x_3\circ (x_2\circ x_1')'=\\
(y_1\succeq y_2)\succeq y_3\otimes(x_1'' x_2`x_3`+x_1'x_2'`x_3`)-y_3\succeq(y_1\succeq y_2)\otimes (x_3'x_1'`x_2`+x_3'x_1'x_2``)\\
+(y_1\succeq y_2)\preceq y_3\otimes x_1' x_2`x_3'`-y_3\preceq(y_1\succeq y_2)\otimes(x_3x_1''` x_2`+x_3x_1''x_2``+x_3x_1'`x_2'`+x_3x_1'x_2'``)\\
-(y_2\succeq y_1)\succeq y_3\otimes(x_2'' x_1`x_3`+x_2'x_1'`x_3`)+y_3\succeq(y_2\succeq y_1)\otimes (x_3'x_2'`x_1`+x_3'x_2'x_1``)\\
-(y_2\succeq y_1)\preceq y_3\otimes x_2' x_1`x_3'`+y_3\preceq(y_2\succeq y_1)\otimes(x_3x_2''` x_1`+x_3x_2''x_1``+x_3x_2'`x_1'`+x_3x_2'x_1'``)
\end{multline*}
\vspace*{-\baselineskip}
\vspace*{-\baselineskip}
\begin{multline*}
+(y_1\preceq y_2)\succeq y_3\otimes(x_1'x_2'`x_3`+x_1 x_2''`x_3`) 
-y_3\succeq(y_1\preceq y_2)\otimes(x_3'x_1`x_2'`+x_3'x_1x_2'``)\\
+(y_1\preceq y_2)\preceq y_3\otimes x_1x_2'`x_3'`-
y_3\preceq(y_1\preceq y_2)\otimes (x_3x_1'`x_2'`+x_3x_1'x_2'``+x_3x_1`x_2''`+x_3x_1x_2''``)\\
-(y_2\preceq y_1)\succeq y_3\otimes(x_2'x_1'`x_3`+x_2 x_1''`x_3`) 
+y_3\succeq(y_2\preceq y_1)\otimes(x_3'x_2`x_1'`+x_3'x_2x_1'``)\\
-(y_2\preceq y_1)\preceq y_3\otimes x_2x_1'`x_3'`+
y_3\preceq(y_2\preceq y_1)\otimes (x_3x_2'`x_1'`+x_3x_2'x_1'``+x_3x_2`x_1''`+x_3x_2x_1''``),
\end{multline*}
where $x\succ y=x'\circ y$, $x\prec y=x\circ y'$ and $x\circ y=xy`$. 
Analogically,
\begin{multline*}
[[y_2\otimes x_2,y_3\otimes x_3],y_1\otimes x_1]=\\
(y_2\succeq y_3)\succeq y_1\otimes(x_2'' x_3`x_1`+x_2'x_3'`x_1`)-y_1\succeq(y_2\succeq y_3)\otimes (x_1'x_2'`x_3`+x_1'x_2'x_3``)\\
+(y_2\succeq y_3)\preceq y_1\otimes x_2' x_3`x_1'`-y_1\preceq(y_2\succeq y_3)\otimes(x_1x_2''` x_3`+x_1x_2''x_3``+x_1x_2'`x_3'`+x_1x_2'x_3'``)\\
-(y_3\succeq y_2)\succeq y_1\otimes(x_3'' x_2`x_1`+x_3'x_2'`x_1`)+y_1\succeq(y_3\succeq y_2)\otimes (x_1'x_3'`x_2`+x_1'x_3'x_2``)\\
-(y_3\succeq y_2)\preceq y_1\otimes x_3' x_2`x_1'`+y_1\preceq(y_3\succeq y_2)\otimes(x_1x_3''` x_2`+x_1x_3''x_2``+x_1x_3'`x_2'`+x_1x_3'x_2'``)\\
+(y_2\preceq y_3)\succeq y_1\otimes(x_2'x_3'`x_1`+x_2 x_3''`x_1`) 
-y_1\succeq(y_2\preceq y_3)\otimes(x_1'x_2`x_3'`+x_1'x_2x_3'``)\\
+(y_2\preceq y_3)\preceq y_1\otimes x_2x_3'`x_1'`-
y_1\preceq(y_2\preceq y_3)\otimes (x_1x_2'`x_3'`+x_1x_2'x_3'``+x_1x_2`x_3''`+x_1x_2x_3''``)\\
-(y_3\preceq y_2)\succeq y_1\otimes(x_3'x_2'`x_1`+x_3 x_2''`x_1`) 
+y_1\succeq(y_3\preceq y_2)\otimes(x_1'x_3`x_2'`+x_1'x_3x_2'``)\\
-(y_3\preceq y_2)\preceq y_1\otimes x_3x_2'`x_1'`+
y_1\preceq(y_3\preceq y_2)\otimes (x_1x_3'`x_2'`+x_1x_3'x_2'``+x_1x_3`x_2''`+x_1x_3x_2''``)
\end{multline*}
and
\begin{multline*}
[[y_3\otimes x_3,y_1\otimes x_1],y_2\otimes x_2]=\\
(y_3\succeq y_1)\succeq y_2\otimes(x_3'' x_1`x_2`+x_3'x_1'`x_2`)-y_2\succeq(y_3\succeq y_1)\otimes (x_2'x_3'`x_1`+x_2'x_3'x_1``)\\
+(y_3\succeq y_1)\preceq y_2\otimes x_3' x_1`x_2'`-y_2\preceq(y_3\succeq y_1)\otimes(x_2x_3''` x_1`+x_2x_3''x_1``+x_2x_3'`x_1'`+x_2x_3'x_1'``)\\
-(y_1\succeq y_3)\succeq y_2\otimes(x_1'' x_3`x_2`+x_1'x_3'`x_2`)+y_2\succeq(y_1\succeq y_3)\otimes (x_2'x_1'`x_3`+x_2'x_1'x_3``)\\
-(y_1\succeq y_3)\preceq y_2\otimes x_1' x_3`x_2'`+y_2\preceq(y_1\succeq y_3)\otimes(x_2x_1''` x_3`+x_2x_1''x_3``+x_2x_1'`x_3'`+x_2x_1'x_3'``)\\
+(y_3\preceq y_1)\succeq y_2\otimes(x_3'x_1'`x_2`+x_3 x_1''`x_2`) 
-y_2\succeq(y_3\preceq y_1)\otimes(x_2'x_3`x_1'`+x_2'x_3x_1'``)\\
+(y_3\preceq y_1)\preceq y_2\otimes x_3x_1'`x_2'`-
y_2\preceq(y_3\preceq y_1)\otimes (x_2x_3'`x_1'`+x_2x_3'x_1'``+x_2x_3`x_1''`+x_2x_3x_1''``)\\
-(y_1\preceq y_3)\succeq y_2\otimes(x_1'x_3'`x_2`+x_1 x_3''`x_2`) 
+y_2\succeq(y_1\preceq y_3)\otimes(x_2'x_1`x_3'`+x_2'x_1x_3'``)\\
-(y_1\preceq y_3)\preceq y_2\otimes x_1x_3'`x_2'`+
y_2\preceq(y_1\preceq y_3)\otimes (x_2x_1'`x_3'`+x_2x_1'x_3'``+x_2x_1`x_3''`+x_2x_1x_3''``).
\end{multline*}








For simplicity, in $\Der\Nov^!$ instead of the operations $\preceq$ and $\succeq$ we write the operations $\prec$ and $\succ$, respectively.
Finally, we obtain the following result:
\begin{proposition}
An operad $\Der\Nov^!$ is defined by the following identities:
\begin{equation}\label{id11}
a\prec(b\prec c)+(b\prec c)\succ a=0,    
\end{equation}
\begin{equation}\label{id12}
a\prec(b\succ c)=0,  
\end{equation}
\begin{equation}\label{id13}
(a\prec b)\prec c-(c\prec b)\succ a=(a\prec c)\prec b-(b\prec c)\succ a,   
\end{equation}
\begin{equation}\label{id14}
(a\succ b)\prec c+c\succ(a\succ b)-(a\succ c)\succ b=0,
\end{equation}
\begin{equation}\label{id15}
a\succ(b\prec c)+(b\prec c)\succ a=0,
\end{equation}
\begin{equation}\label{id16}
a\succ(b\succ c)=b\succ(a\succ c),
\end{equation}
\begin{equation}\label{id17}
(a\prec b)\succ c=(c\prec b)\succ a
\end{equation}
and
\begin{equation}\label{id18}
(a\succ b)\succ c=(a\succ c)\succ b.
\end{equation}    
\end{proposition}

Let us denote by $\Der\Nov^!\<X\>$ a free algebra generated by the countable set $X$ defined by the identities of the operad $\Der\Nov^!$. Indeed, every monomial in $\Der\Nov^!\<X\>$ can be expressed as a sum of pure monomials involving only the operations $\prec$ or $\succ$. For monomials of the form
\[
(a\prec b)\succ c,\; (a\succ b)\prec c,\; a\prec (b\succ c)\; \textrm{and}\; a\succ (b\prec c),
\] 
we apply identities (\ref{id11}), (\ref{id14}), (\ref{id12}) and (\ref{id15}), respectively. Applying these identities to (\ref{id13}) and (\ref{id17}), we obtain the following result:
\begin{proposition}
\[
\Der\Nov^!\<X\>=\Der\Nov^!_{\prec}\<X\>\oplus\Der\Nov^!_{\succ}\<X\>,
\]
as a vector space, where $\Der\Nov^!_{\prec}\<X\>$ and $\Der\Nov^!_{\succ}\<X\>$ are the subspaces of $\Der\Nov^!\<X\>$ generated by monomials containing only operations $\prec$ and $\succ$, respectively. In particular, $\Der\Nov^!_{\prec}\<X\>$ is a Novikov algebra with some additional identities of a higher degree.
Analogically, $\Der\Nov^!_{\succ}\<X\>$ is a bicommutative algebra with some additional identities of a higher degree.
\end{proposition}
We aim to describe a complete list of identities that appear in $\Der\Nov^!_{\prec}\<X\>$ and $\Der\Nov^!_{\succ}\<X\>$.

\begin{theorem}\label{th1}
The algebra $\Der\Nov^!_{\prec}\<X\>$ coincides with a free Novikov algebra with the following additional identities:
\begin{equation}\label{Nov1}
(a\prec(b\prec c))\prec d=a\prec(b\prec (c\prec d))=0,
\end{equation}
\begin{equation}\label{Nov2}
a\prec((b\prec c)\prec d)=a\prec((b\prec d)\prec c),
\end{equation}
and
\begin{equation}\label{Nov3}
((a\prec b)\prec c)\prec d=((a\prec c)\prec b)\prec d.
\end{equation}
\end{theorem}
\begin{proof}
To prove the result, we consider the identities (\ref{id11})-(\ref{id18}) as a rewriting system and compute the Gr\"obner basis for the corresponding operad. Let us define the order on monomials as follows:
\begin{itemize}
    \item A monomial involving various types of operations is always greater than a monomial that contains only one type of operation.
    \item Monomials involving various types of operations are ordered in lexicographic order, where the operation $\succ$ is greater than the operation $\prec$.
    \item Monomials that contain only one type of operation are ordered in lexicographic order.
\end{itemize}

First, we show that identities (\ref{Nov1})-(\ref{Nov3}) hold in algebra $\Der\Nov^!\<X\>$. From the composition of (\ref{id11}) and (\ref{id12}), we obtain
\[
a\prec((c\prec d)\succ b)=a\prec(b\prec(c\prec d))=0.
\]
The identity
\[
a\succ(b\prec c)=a\prec(b\prec c)
\]
can be obtained from (\ref{id11}) and (\ref{id15}), and the identity
\[
(b\prec a)\succ c=(c\prec a)\prec b-(c\prec b)\prec a-a\prec(c\prec b)
\]
can be obtained from (\ref{id11}) and (\ref{id13}). From the composition of the last two given identities, we have
\begin{multline*}
(b\prec a)\succ(c\prec d)=((c\prec d)\prec a)\prec b-((c\prec d)\prec b)\prec a-a\prec ((c\prec d)\prec b)= \\
(c\prec d)\prec (a\prec b)+((c\prec d)\prec b)\prec a-(c\prec d)\prec (b\prec a)-((c\prec d)\prec b)\prec a-a\prec ((c\prec d)\prec b)=\\
-(c\prec d)\prec (b\prec a)
\end{multline*}
and
\begin{multline*}
(b\prec a)\succ(c\prec d)=(b\prec a)\prec(c\prec d)=b\prec(a\prec(c\prec d))+(b\prec (c\prec d))\prec a\\
-b\prec((c\prec d)\prec a).
\end{multline*}
The equality gives
\[
b\prec(a\prec(c\prec d))+(b\prec (c\prec d))\prec a=(b\prec (c\prec d))\prec a=0.
\]
The identities (\ref{Nov2}) and (\ref{Nov3}) can be easily proved using (\ref{Nov1}) and the right-symmetric identity.

To prove the result, we must show that all compositions of rewriting rules are trivial, see \cite{bremner-dotsenko}. Using the computer software \cite{DotsHij}, we find that at degree 4, all compositions involving various operations are either trivial or equal to zero. For compositions of higher degrees, only rewriting rules involving a single operation remain.
\end{proof}

\begin{theorem}
The algebra $\Der\Nov^!_{\succ}\<X\>$ coincides with a free bicommutative algebra with the following additional identities:
\begin{equation}\label{bicom1}
((a\succ b)\succ c)\succ d=d\succ(c\succ(b\succ a))=c\succ(b\succ(a\succ d)).  
\end{equation}
\begin{equation}\label{bicom2}
((a\succ b)\succ c)\succ d=c\succ((d\succ a)\succ b)=b\succ((d\succ a)\succ c)=b\succ((c\succ a)\succ d).  
\end{equation}
\end{theorem}
\begin{proof}
We prove the result analogically as proved Theorem \ref{th1}. First, we show that the identities (\ref{bicom1}) and (\ref{bicom2}) hold in the algebra $\Der\Nov^!\<X\>$.
From (\ref{id11}) and (\ref{id15}), we have
\[
b\succ(a\prec c)=a\succ(b\prec c).
\]
Calculating the composition of the obtained identity and (\ref{id14}), one can obtain 
\[
(b\succ d)\succ (a\prec c)=a\succ((b\succ d)\prec c)=^{(\ref{id14})} -a\succ(c\succ (b\succ d))+a\succ((b\succ c)\succ d)
\]
and
\begin{multline*}
(b\succ d)\succ (a\prec c)=(b\succ d)\prec (a\prec c)+(a\prec c)\succ (b\succ d)=^{(\ref{id14})}\\
-(a\prec c)\succ (b\succ d)+(b\succ(a\prec c))\succ d+(a\prec c)\succ (b\succ d)=(b\succ(a\prec c))\succ d=^{(\ref{id15})}\\
((a\prec c)\succ b)\succ d=^{(\ref{id11})} (b\prec (a\prec c))\succ d=^{(\ref{id11})} d\prec(b\prec (a\prec c))=^{(\ref{Nov1})}0.
\end{multline*}
The equality gives
\begin{equation}\label{help1}
a\succ(c\succ (b\succ d))=a\succ((b\succ c)\succ d).    
\end{equation}

The identities (\ref{id11}) and (\ref{id15}) gives
\[
a\succ(b\prec c)=a\prec(b\prec c).
\]
Calculating the composition of this identity and (\ref{id14}), we obtain
\[
a\succ((b\succ c)\prec d)=a\prec((b\succ c)\prec d)=^{(\ref{id14})} -a\prec(d\succ(b\succ c))+a\prec( (b\succ d)\succ c)=^{(\ref{id12})} 0
\]
and
\begin{multline*}
a\succ((b\succ c)\prec d)=a\succ((b\succ d)\prec c)-a\succ(b\succ(d\succ c))+a\succ(b\succ(c\succ d))=^{(\ref{id14})} \\
-a\succ(c\succ(b\succ d))+a\succ((b\succ c)\succ d)-a\succ(b\succ(d\succ c))+a\succ(b\succ(c\succ d))=^{(\ref{help1})} \\
-a\succ(b\succ(d\succ c))+a\succ(b\succ(c\succ d)).
\end{multline*}
The equality gives
\begin{equation}\label{help2}
a\succ(b\succ(d\succ c))=a\succ(b\succ(c\succ d)).    
\end{equation}

The last necessary identity can be obtained from the composition of (\ref{id14}) and (\ref{id15}). So, we have
\[
((a\succ d)\prec b)\succ c=c\succ((a\succ d)\prec b)=^{(\ref{id14})} -c\succ(b\succ(a\succ d))+c\succ((a\succ b)\succ d)=^{(\ref{help1})} 0
\]
and
\[
((a\succ d)\prec b)\succ c=((a\succ d)\succ b)\succ c-(a\succ (b\succ d))\succ c.
\]
The equality gives
\begin{equation}\label{help3}
((a\succ d)\succ b)\succ c=(a\succ (b\succ d))\succ c.   
\end{equation}
The identities (\ref{bicom1}) and (\ref{bicom2}) can be easily proved by using (\ref{help1}), (\ref{help2}) and (\ref{help3}).

To prove the result, as before, we have to show that all compositions of rewriting rules are trivial. The calculations are straightforward and all these long calculations can be done using the computer software \cite{DotsHij}.

\end{proof}

\section{A basis of algebra $\Der\Nov^!\<X\>$}

To construct a basis of the algebra $\Der\Nov^!\<X\>$ it suffices to construct a basis for the algebras $\Der\Nov^!_{\succ}\<X\>$ and $\Der\Nov^!_{\prec}\<X\>$ separately. 

Let us define a set $\mathcal{N}_i$ as follows:
\[
\mathcal{N}_1=\{x_i\},\;\mathcal{N}_2=\{x_ix_j\},\;\mathcal{N}_3=\{(x_{k_1}x_{k_2})x_{k_3},x_{l_1}(x_{l_2}x_{l_3})\}
\]
and
\[
\mathcal{N}_n=\{((\cdots(x_{m_1}x_{m_{2}})\cdots)x_{m_{n-1}})x_{m_n}, x_{r_1}((\cdots(x_{r_2}x_{r_{3}})\cdots)x_{r_n})\},
\]
where $k_2\leq k_3$, $l_1\geq l_2$, $n\geq 4$, $m_n\geq \ldots \geq m_2$ and $r_2\geq r_1$, $r_n\geq\ldots\geq r_3$. Also, we set
\[
\mathcal{N}=\bigcup_{i}\mathcal{N}_i
\]
\begin{theorem}
Let $\Nov_s\<X\>$ be a free Novikov algebra with identities (\ref{Nov1}) and (\ref{Nov2}). The set $\mathcal{N}$ is the basis of the algebra $\Nov_s\<X\>$.
\end{theorem}
\begin{proof}
For monomials up to degree $4$, the result is obvious. For monomials bigger than degree $4$, we first show that any monomial from $\Nov_s\<X\>$ can be expressed as a sum of monomials of the form
\[
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-1}})x_{i_n}\;\textrm{and}\; x_{i_1}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_n}).
\]
Let us denote by $A_i$ an arbitrary monomial of degree $i$. Starting from degree $5$, we prove it by induction on the length of monomials. Consider $4$ cases. By the inductive hypothesis, we have

Case 1: $(A_{n-1})(A_1)$. If $(A_{n-1})=(\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-1}}$ then there is nothing to do. Otherwise, $(A_{n-1})(A_1)=^{(\ref{Nov1})}0$.

Case 2.1: By left-commutative identity,
\[
(A_{n-2})(A_2)=((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})(x_{i_{n-1}}x_{i_{n}})=x_{i_{n-1}}(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})x_{i_{n}}).
\]

Case 2.2: By left-commutative identity,
\begin{multline*}
(A_{n-2})(A_2)=(x_{i_1}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_{n-2}}))(x_{i_{n-1}}x_{i_{n}})=\\
x_{i_{n-1}}((x_{i_1}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_{n-2}}))x_{i_{n}})=^{(\ref{Nov1})}0.    
\end{multline*}

Case 3.1: By right-symmetric identity:
\begin{multline*}
(A_{n-3})(A_3)=((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})((x_{i_{n-2}}x_{i_{n-1}})x_{i_n})=
(\cdots(x_{i_1}x_{i_{2}})\cdots)(x_{i_{n-3}}((x_{i_{n-2}}x_{i_{n-1}})x_{i_n}))\\
+((\cdots(x_{i_1}x_{i_{2}})\cdots)((x_{i_{n-2}}x_{i_{n-1}})x_{i_n}))x_{i_{n-3}}-(\cdots(x_{i_1}x_{i_{2}})\cdots)(((x_{i_{n-2}}x_{i_{n-1}})x_{i_n})x_{i_{n-3}})=^{(\ref{Nov1})}\\
-((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-4}})(((x_{i_{n-2}}x_{i_{n-1}})x_{i_n})x_{i_{n-3}}).
\end{multline*}
Applying right-symmetric identity and (\ref{Nov1}) several times as given above we obtain
\[
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})((x_{i_{n-2}}x_{i_{n-1}})x_{i_n})=(-1)^{n-4} x_{i_1}((\cdots(((x_{i_{n-2}}x_{i_{n-1}})x_{i_n})x_{i_{n-3}})\cdots)x_{i_2}).
\]

Case 3.2: By (\ref{Nov1}),
\begin{multline*}
(A_{n-3})(A_3)=((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})(x_{i_{n-2}}(x_{i_{n-1}}x_{i_n}))=\\
(x_{i_{n-4}}((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}}))((x_{i_{n-2}}x_{i_{n-1}})x_{i_n})=\\
(x_{i_{n-4}}((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}}))(x_{i_{n-2}}(x_{i_{n-1}}x_{i_n}))=0.
\end{multline*}

Case 4.1: As given in case 3.1 we use right-symmetric identity and (\ref{Nov1}), and obtain 
\begin{multline*}
(A_{n-k})(A_{k})=((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-k}})((\cdots(x_{i_{n-k+1}}x_{i_{n-k+2}})\cdots)x_{i_{n}})=\\
(-1)^{n-k-1} x_{i_1}((\cdots(((\cdots(x_{i_{n-k+1}}x_{i_{n-k+2}})\cdots)x_{i_{n}})x_{i_{n-k}})\cdots)x_{i_2}),
\end{multline*}
where $k>3$.

Case 4.2: By (\ref{Nov1}),
\begin{multline*}
(A_{n-k})(A_{k})=(x_{i_{1}}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_{n-k}}))((\cdots(x_{i_{n-k+1}}x_{i_{n-k+2}})\cdots)x_{i_{n}})=\\
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-k}})(x_{i_{n-k+1}}((\cdots(x_{i_{n-k+2}}x_{i_{n-k+3}})\cdots)x_{i_{n}}))=\\
(x_{i_{1}}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_{n-k}}))(x_{i_{n-k+1}}((\cdots(x_{i_{n-k+2}}x_{i_{n-k+3}})\cdots)x_{i_{n}}))=0.
\end{multline*}

Now, let us show that all monomials of the form
\[
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-1}})x_{i_n}\;\textrm{and}\; x_{i_1}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_n}).
\]
can be written as a sum monomials from $\mathcal{N}$.
By right-symmetric identity and (\ref{Nov1}),
\begin{multline*}
(\cdots(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_t})x_{i_{t+1}})\cdots)x_{i_n}=(\cdots((\cdots(x_{i_1}x_{i_{2}})\cdots)(x_{i_t}x_{i_{t+1}}))\cdots)x_{i_n}\\
+(\cdots(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{t+1}})x_{i_{t}})\cdots)x_{i_n}-(\cdots((\cdots(x_{i_1}x_{i_{2}})\cdots)(x_{i_{t+1}}x_{i_{t}}))\cdots)x_{i_n}=\\
(\cdots(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{t+1}})x_{i_{t}})\cdots)x_{i_n},
\end{multline*}
that is, the generators $x_{i_t}$ and $x_{i_{t+1}}$ can be ordered, where $1<t<n-1$. For $t=n-1$, using the case 3.1, we have
\begin{multline*}
((((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})x_{i_{n-1}})x_{i_n}=((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})(x_{i_{n-1}}x_{i_n})\\
+(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})x_{i_{n}})x_{i_{n-1}}-(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})(x_{i_{n}}x_{i_{n-1}})=\\
(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})x_{i_{n}})x_{i_{n-1}}+(-1)^{n-3}
x_{i_1}((\cdots((x_{i_{n-1}}x_{i_n})x_{i_{n-2}})\cdots)x_{i_{2}})\\
-(-1)^{n-3}
x_{i_1}((\cdots((x_{i_{n}}x_{i_{n-1}})x_{i_{n-2}})\cdots)x_{i_{2}}),
\end{multline*}
that is, the generators $x_{i_{n-1}}$ and $x_{i_{n}}$ can be ordered by modulo of monomials of the form $x_{i_1}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_n})$. In an analogical way, it can be shown that the generators $x_{i_{3}},x_{i_{4}},\ldots,x_{i_{n}}$ in $x_{i_1}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_n})$ can be ordered.

Let us apply left-commutative and right-symmetric identities on monomial 
\[
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})(x_{i_{n-1}}x_{i_n})
\] in two different ways as follows:
\[
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})(x_{i_{n-1}}x_{i_n})=x_{i_{n-1}}(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})x_{i_n})
\]
and
\begin{multline*}
(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})x_{i_{n-2}})(x_{i_{n-1}}x_{i_n})=((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})(x_{i_{n-2}}(x_{i_{n-1}}x_{i_n}))\\
+(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})(x_{i_{n-1}}x_{i_n}))x_{i_{n-2}}-((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})((x_{i_{n-1}}x_{i_n})x_{i_{n-2}})=^{(\ref{Nov1})}\\
-((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-3}})((x_{i_{n-1}}x_{i_n})x_{i_{n-2}})=\cdots=\\
(-1)^{n-3}
x_{i_1}((\cdots(((x_{i_{n-1}}x_{i_n})x_{i_{n-2}})x_{i_{n-3}})\cdots)x_{i_{2}}).
\end{multline*}
So, we obtain 
\[
x_{i_1}((\cdots(((x_{i_{n-1}}x_{i_n})x_{i_{n-2}})x_{i_{n-3}})\cdots)x_{i_{2}})=(-1)^{n-3}x_{i_{n-1}}(((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n-2}})x_{i_n}),
\]
that is, the generators $x_{i_1}$ and $x_{i_{n-1}}$ can be ordered.

Consider an algebra $A\<X\>$ with a basis $\mathcal{N}$ and multiplication on basis monomials is defined as follows:

\begin{itemize}
    \item Up to degree 3, we define multiplication in $A\<X\>$ that is consistent with left-commutative and right-symmetric identities.
    \item $((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_n})((\cdots(x_{j_1}x_{j_{2}})\cdots)x_{j_m})=x_{k_1}((\cdots(x_{k_2}x_{k_{3}})\cdots)x_{k_{n+m}})$, where $k_1\leq k_2$, $k_3\leq\ldots\leq k_{n+m}$ and $\{k_1,k_2\}=\{i_1,j_1
    \}$, $\{k_3,\ldots,k_{n+m}\}=\{i_2,\ldots,i_n,j_2,\ldots,j_m\}$.
    \item $(x_{i_{1}}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_{n}}))((\cdots(x_{j_{1}}x_{j_{2}})\cdots)x_{j_{m}})=\\
((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_{n}})(x_{j_{1}}((\cdots(x_{j_{2}}x_{j_{3}})\cdots)x_{j_{m}}))=\\
(x_{i_{1}}((\cdots(x_{i_2}x_{i_{3}})\cdots)x_{i_{n}}))(x_{j_{1}}((\cdots(x_{j_{2}}x_{j_{3}})\cdots)x_{j_{m}}))=0.$
\end{itemize}
By straightforward calculation, one can check that an algebra $A\<X\>$ satisfies the defining identities of algebra $\Nov_s\<X\>$. It remains to note that $A\<X\>\cong \Nov_s\<X\>$.
\end{proof}

Let us define a set $\mathcal{B}_i$ as follows:
\[
\mathcal{B}_1=\{x_i\},\;\mathcal{B}_2=\{x_ix_j\},\;\mathcal{B}_3=\{(x_{k_1}x_{k_2})x_{k_3},x_{l_1}(x_{l_2}x_{l_3})\}
\]
and
\[
\mathcal{B}_n=\{((\cdots(x_{m_1}x_{m_{2}})\cdots)x_{m_{n-1}})x_{m_n}\},
\]
where $k_2\leq k_3$, $l_1\geq l_2$, $n\geq 4$ and $m_n\geq \ldots \geq m_2\geq m_1$. Also, we set
\[
\mathcal{B}=\bigcup_{i}\mathcal{B}_i
\]
\begin{theorem}
Let $\Bi\Com_s\<X\>$ be a free bicommutative algebra with identities (\ref{bicom1}) and (\ref{bicom2}). The set $\mathcal{B}$ is the basis of the algebra $\Bi\Com_s\<X\>$.
\end{theorem}
\begin{proof}
For monomials up to degree $4$, the result is obvious.  For monomials of degree greater than $4$, we first show that any monomial in $\Bi\Com_s\<X\>$ can be written as a sum of left-normed monomials.  Let us denote by $A_i$ an arbitrary monomial of degree $i$. Starting from degree $5$, we prove it by induction on the length of monomials. Consider $4$ cases. By the inductive hypothesis, we have

Case 1: 
\begin{multline*}
(A_1)(A_{n-1})=x_{i_1}((\cdots(x_{i_2}x_{i_3})\cdots)x_{i_{n-1}})x_{i_n}=^{(\ref{bicom2})}((x_{i_{n-1}}x_{i_{n}})x_{i_1})((\cdots(x_{i_2}x_{i_3})\cdots)x_{i_{n-2}})= \\
((x_{i_{n-1}}x_{i_{n}})((\cdots(x_{i_2}x_{i_3})\cdots)x_{i_{n-2}}))x_{i_1}=((\cdots(x_{j_1}x_{j_2})\cdots)x_{j_{n-1}})x_{i_1}. 
\end{multline*}

Case 2:
\begin{multline*}
(A_2)(A_{n-2})=(x_{i_1}x_{i_2})((\cdots(x_{i_3}x_{i_4})\cdots)x_{i_{n-1}})x_{i_n})=(x_{i_1}((\cdots(x_{i_3}x_{i_4})\cdots)x_{i_{n-1}})x_{i_n}))x_{i_2}=\\((\cdots(x_{j_1}x_{j_2})\cdots)x_{j_{n-1}})x_{i_2}. 
\end{multline*}

Case 3.1:
\begin{multline*}
(A_3)(A_{n-3})=((x_{i_1}x_{i_2})x_{i_3})((\cdots(x_{i_4}x_{i_5})\cdots)x_{i_n})=((x_{i_1}x_{i_2})((\cdots(x_{i_4}x_{i_5})\cdots)x_{i_n}))x_{i_3}=\\((\cdots(x_{j_1}x_{j_2})\cdots)x_{j_{n-1}})x_{i_3}. 
\end{multline*}

Case 3.2:
\begin{multline*}
(A_3)(A_{n-3})=(x_{i_1}(x_{i_2}x_{i_3}))((\cdots(x_{i_4}x_{i_5})\cdots)x_{i_n})=(x_{i_1}((\cdots(x_{i_4}x_{i_5})\cdots)x_{i_n}))(x_{i_2}x_{i_3})=\\
x_{i_2}((x_{i_1}((\cdots(x_{i_4}x_{i_5})\cdots)x_{i_n}))x_{i_3})
\end{multline*}
which correspond to the case 1.

Case 4:
\begin{multline*}
(A_k)(A_{n-k})=((\cdots(x_{i_1}x_{i_2})\cdots)x_{i_k})(A_{n-k})=((\cdots(x_{i_1}x_{i_2})\cdots)(A_{n-k}))x_{i_k}=\\
((\cdots(x_{j_1}x_{j_2})\cdots)x_{j_{n-1}})x_{i_k}.
\end{multline*}
For an arbitrary monomial of the form
\[
((\cdots(x_{m_1}x_{m_{2}})\cdots)x_{m_{n-1}})x_{m_n}
\]
the generators $x_{m_{2}},\ldots,x_{m_{n}}$ can be ordered by the right-commutative identity. By (\ref{bicom2}), we have
\[
((ab)c)d=c((da)b)=c((db)a)=((ba)c)d,
\]
that is, the generators $x_{m_{1}}$ and $x_{m_{2}}$ can also be ordered.

Consider an algebra $A\<X\>$ with a basis $\mathcal{B}$ and multiplication on basis monomials is defined as follows:

\begin{itemize}
    \item Up to degree 3, we define multiplication in $A\<X\>$ that is consistent with right-commutative and left-commutative identities.
    \item $((\cdots(x_{i_1}x_{i_{2}})\cdots)x_{i_n})((\cdots(x_{j_1}x_{j_{2}})\cdots)x_{j_m})=(\cdots(x_{k_1}x_{k_{2}})\cdots)x_{k_{n+m}}$, where $k_1\leq\ldots\leq k_{n+m}$ and $\{k_1,\ldots,k_{n+m}\}=\{i_1,\ldots,i_n,j_1,\ldots,j_m\}$.
\end{itemize}
By straightforward calculation, one can check that an algebra $A\<X\>$ satisfies the defining identities of algebra $\Bi\Com_s\<X\>$. It remains to note that $A\<X\>\cong \Bi\Com_s\<X\>$.
\end{proof}

\begin{proposition}
An operad $\Bi\Com$ satisfies the Dong property in the sense of \cite{Dongprop}.
\end{proposition}
\begin{proof}
The Lie-admissibility condition for $S\otimes U$ gives the defining identities of the operad $\Bi\Com^!$, where $U$ is a $\Bi\Com$ algebra.
Then
\begin{multline*}
[[a\otimes u,b\otimes v],c\otimes w]=(ab)c\otimes (uv)w-(ba)c\otimes (vu)w-
c(ab)\otimes w(uv)+c(ba)\otimes w(vu),
\end{multline*}
\begin{multline*}
[[b\otimes v,c\otimes w],a\otimes u]=(bc)a\otimes (vw)u-(cb)a\otimes (wv)u-
a(bc)\otimes u(vw)+a(cb)\otimes u(wv),
\end{multline*}
and
\begin{multline*}
[[c\otimes w,a\otimes u],b\otimes v]=(ca)b\otimes (wu)v-(ac)b\otimes (uw)v-
b(ca)\otimes v(wu)+b(ac)\otimes v(uw).
\end{multline*}
Calculating the sum and collecting the same basis monomials on the right side of the tensors, we obtain
\[
(ab)c=(ac)b\;\;\textrm{and}\;\; a(bc)=b(ac).
\]
So, an operad $\Bi\Com$ is self-dual and the monomials $(ab)c$, $(ba)c$, $c(ab)$ and $c(ba)$ are linearly independent in $\Bi\Com^!\<X\>$. By the criterion given in \cite{Dongprop}, the operad $\Bi\Com$ satisfies the Dong property.
\end{proof}

\begin{corollary}
An operad $\Der\Nov^!$ satisfies the Dong property.
\end{corollary}
\begin{proof}
Since the quadratic operads $\Nov$ and $\Bi\Com$ satisfy the Dong property and
\[
\Der\Nov^!(3)=\Nov(3)\oplus\Bi\Com(3),
\]
we obtain the result.
\end{proof}

In addition, we have
\[
\dim(\Der\Nov^!(1))=1,\dim(\Der\Nov^!(2))=2,\dim(\Der\Nov^!(3))=12
\]
and starting from degree $4$,
\[
\dim(\Der\Nov^!(n))=n(n+1)/2+1.
\]

\end{document}